# Characterizing Kadison–Schwarz maps on $M_{3}$

### v2

@@ Abstract
Kadison–Schwarz (KS) maps form a natural class of positive linear maps lying between positivity and complete positivity. Despite their relevance in quantum dynamics and operator algebras, a detailed analytic characterization of KS maps is still largely lacking. In this work we analyze unital linear maps on $M_{3}$ using the Bloch–Gell–Mann representation. Exploiting unitary equivalence and structural properties of the $\mathfrak{su}(3)$ algebra, we derive explicit analytic conditions ensuring the Kadison–Schwarz property. Our approach clarifies the relation between KS maps and completely positive maps on $M_{3}$.
@@ I Introduction
Positive linear maps between matrix algebras play a fundamental role in quantum information theory, operator algebras, and the theory of open quantum systems. Among various classes of positive maps, completely positive (CP) maps occupy a distinguished position due to their physical interpretation as quantum channels [Choi1975]. However, in several contexts—notably in the analysis of quantum dynamical maps, divisibility, and non-Markovian dynamics—one encounters broader classes of positive maps for which complete positivity is not required, see, e.g., Refs. [Chruscinski2010, Chruscinski2017].
@@
Several general features of KS maps are known. In particular, unital positive maps satisfy the Kadison–Schwarz inequality when restricted to normal operators, whereas explicit counterexamples show that unitality and positivity alone do not guarantee the KS property. Beyond such general observations, however, explicit analytic characterizations of KS maps remain scarce and are typically restricted to low-dimensional or highly symmetric settings.
The aim of the present work is to analyze Kadison–Schwarz maps within the Bloch–Gell–Mann framework. This representation provides a natural parametrization of linear maps on $M_{d}$ adapted to the $\mathfrak{su}(d)$ structure and allows one to systematically exploit unitary equivalence and canonical decompositions of the associated Bloch matrix. While the framework itself applies to arbitrary dimension $d$, the derivation of explicit analytic conditions depends sensitively on the underlying algebraic structure.
In this work we derive explicit analytic sufficient conditions for the Kadison–Schwarz property in the case of the matrix algebra $M_{3}$. The special role of $M_{3}$ stems from the tractable structure of $\mathfrak{su}(3)$, which permits nontrivial analytic bounds without resorting to numerical optimization. Our results clarify the relation between Kadison–Schwarz maps and completely positive maps on $M_{3}$ and identify structural features of the Bloch–Gell–Mann representation that are responsible for the KS property. To the best of our knowledge, explicit analytic sufficient conditions for the Kadison–Schwarz property of unital maps on $M_{3}$, expressed directly in terms of Bloch (Gell–Mann) parameters, have not been systematically discussed in the literature beyond the qubit case ($d=2$) or highly symmetric subclasses.
The paper is organized as follows. In Sec. II we recall the definition of Kadison–Schwarz maps and review their basic properties. In Sec. III we introduce the Bloch–Gell–Mann representation of unital maps on $M_{d}$. Our main results, specialized to the case of $M_{3}$, are presented in Sec. IV. Technical derivations are collected in the Appendix.
@@ II Kadison–Schwarz maps
which is required to hold for all $X\in M_{d}$. The inequality (1) was originally introduced by Kadison in the context of operator algebras [Kadison1952]. We emphasize that the KS property does not, in general, imply unitality, although the two notions are often considered together.
@@
see, e.g., Refs. [Chruscinski2010, Mukhamedov2011]. Despite this simple ordering, the structure of KS maps is considerably less understood than that of completely positive maps.
@@
Lemma 1 (Kadison inequality).
@@
Lemma 1 implies that unital positive maps are contractions on the space of Hermitian operators. Importantly, this property does not rely on complete positivity and plays a key role in the analysis of KS maps [Kadison1952, Chruscinski2010].
It is worth noting that for unital positive maps the Kadison–Schwarz inequality is automatically satisfied when restricted to normal operators, whereas possible violations of the KS condition necessarily involve non-normal elements [Chruscinski2010, Mukhamedov2011].
@@
Hence the inequality $T(X^{\dagger}X)\geq T(X^{\dagger})T(X)\;$ is violated. This example is well known and discussed in detail in the literature, see, e.g., Refs. [Chruscinski2010, Mukhamedov2011].
@@ III Bloch–Gell–Mann representation of unital maps on $M_{d}$
We emphasize that the Bloch representation of linear maps does not depend on the particular choice of an orthonormal basis of the Lie algebra $\mathfrak{su}(d)$. Different choices, such as the Weyl or generalized Pauli bases, are related by orthogonal transformations in the space of Bloch vectors and therefore lead to equivalent descriptions of unital maps [Kimura2003, BengtssonZyczkowski]. In particular, the Kadison–Schwarz property is invariant under such basis changes. The choice of the Gell–Mann basis is motivated by technical convenience: its generators are Hermitian and possess well-known symmetric and antisymmetric structure constants, which allow for a transparent algebraic treatment of operator products appearing in the Kadison–Schwarz inequality [Georgi1999]. This makes the Gell–Mann representation particularly suitable for deriving analytic bounds.
@@ III.1 Bloch representation on $M_{d}$
where $w_{0}\in\mathbb{R}$ and $w_{k}\in\mathbb{C}$. For Hermitian operators $X$, all coefficients $w_{k}$ are real [Kimura2003, BengtssonZyczkowski].
@@
where $f_{ijk}$ and $d_{ijk}$ are the completely antisymmetric and symmetric structure constants of $\mathfrak{su}(d)$, respectively [Georgi1999, Slansky].
@@ III.2 Unital maps in the Bloch representation
where $w=(w_{1},\dots,w_{d^{2}-1})^{T}$ and $T$ is a real $(d^{2}-1)\times(d^{2}-1)$ matrix, referred to as the Bloch matrix of $\Phi$ [Kimura2003, Ruskai2002]. Unitality of $\Phi$ implies that the identity component $w_{0}$ is preserved, while positivity imposes nontrivial constraints on $T$.
@@
where $O_{1},O_{2}\in SO(d^{2}-1)$ are orthogonal matrices induced by the adjoint action of $SU(d)$ on $\mathfrak{su}(d)$ [BengtssonZyczkowski].
@@ III.3 Canonical form
where $R\in SO(d^{2}-1)$ is orthogonal and $S$ is symmetric and positive semidefinite. Since orthogonal transformations correspond to unitary conjugations, which leave the Kadison–Schwarz property invariant, one may restrict attention to Bloch matrices of the form $T=S$ [Ruskai2002].
Consequently, without loss of generality, unital maps can be represented by symmetric Bloch matrices. In particular, if $S$ is diagonalizable, the map $\Phi$ is unitarily equivalent to a map with diagonal Bloch matrix
@@ IV Main result: Kadison–Schwarz condition on $M_{3}$
Let $\Phi:M_{3}\to M_{3}$ be a unital linear map with diagonal Bloch matrix
@@
in the Gell–Mann representation. Assume that
$\max_{1\leq k\leq 8}|\mu_{k}|\leq 1$ (12)
and that the parameters $\{\mu_{k}\}$ satisfy
$\max_{i,j}|\mu_{i}-\mu_{j}|\;\leq\;\frac{2\sqrt{2}}{3\,C_{3}},$ (13)
where $C_{3}$ is a positive constant depending only on the symmetric structure constants $d_{ijk}$ of $\mathfrak{su}(3)$. Then the map $\Phi$ satisfies the Kadison–Schwarz inequality.
@@
The constant $C_{3}$ arises from norm estimates of expressions involving the symmetric structure constants $d_{ijk}$ of $\mathfrak{su}(3)$. Its explicit value is not essential for the formulation of the result and is discussed in the Appendix. The condition (13) is sufficient but not necessary for the Kadison–Schwarz property.
@@ IV.1 Reduction of the Kadison–Schwarz inequality
A crucial simplification occurs due to the antisymmetric structure constants $f_{ijk}$ of $\mathfrak{su}(3)$. Since the Bloch matrix $T$ is diagonal and the parameters $\mu_{k}$ are real, all terms proportional to $f_{ijk}$ cancel identically in (14). As a consequence, the nontrivial part of the Kadison–Schwarz inequality is governed solely by contributions involving the symmetric structure constants $d_{ijk}$. This mechanism relies essentially on the presence of nonvanishing symmetric structure constants and therefore does not occur in the qubit case $d=2$, where $d_{ijk}$ vanish identically. It is thus specific to $M_{3}$ and higher-dimensional matrix algebras and plays a key role in enabling analytic estimates.
@@ IV.2 Proof of Theorem 1
By the reduction described above, the Kadison–Schwarz condition is equivalent to the positivity of the operator (14). The identity component $\alpha(X)$ is nonnegative under the contraction assumption (12). The remaining traceless part can be estimated using operator norm bounds for expressions involving the symmetric structure constants $d_{ijk}$.
More precisely, one obtains an inequality of the form
@@
where $C_{3}$ depends only on the algebraic structure of $\mathfrak{su}(3)$. The condition (13) guarantees that the traceless contribution in (14) is dominated by the positive identity term. Hence,
@@
for all $X\in M_{3}$, which proves the theorem. A detailed derivation of the bound (15) and an explicit expression for the constant $C_{3}$ are provided in the Appendix.

### v3

# Sufficient conditions for the Kadison–Schwarz property of unital positive maps on $M_{3}$

@@ Abstract
Kadison–Schwarz (KS) maps form a natural class of positive linear maps lying strictly between positivity and complete positivity. Despite their relevance in operator algebras and quantum dynamics, explicit analytic sufficient conditions for the KS property remain scarce beyond low-dimensional or highly symmetric settings. In this work we analyze unital positive linear maps on $M_{3}$ within the Bloch–Gell–Mann representation and derive explicit analytic sufficient conditions ensuring the Kadison–Schwarz property. The approach exploits unitary equivalence together with structural properties of the Lie algebra $\mathfrak{su}(3)$ and does not rely on numerical optimization or semidefinite-programming methods. A key mechanism is the cancellation of contributions associated with antisymmetric structure constants, which reduces the problem to estimates governed solely by the symmetric tensor $d_{ijk}$. The results clarify how the Kadison–Schwarz property can hold under assumptions substantially weaker than complete positivity and yield a structural criterion for the KS property on $M_{3}$ in terms of Bloch parameters.
@@ I Introduction
Positive linear maps between matrix algebras play a fundamental role in quantum information theory, operator algebras, and the theory of open quantum systems. Among various classes of positive maps, completely positive (CP) maps occupy a distinguished position due to their physical interpretation as quantum channels [4, 11, 19]. However, in several contexts—notably in the analysis of quantum dynamical maps, divisibility, and non-Markovian dynamics—one encounters broader classes of positive maps for which complete positivity is not required, see, e.g., Refs. [3, 16, 22, 2, 6].
@@
Several general features of KS maps are known. In particular, unital positive maps satisfy the Kadison–Schwarz inequality when restricted to normal operators, whereas explicit counterexamples show that unitality and positivity alone do not guarantee the KS property. Beyond such general observations, however, explicit analytic sufficient conditions ensuring the KS property remain scarce and are typically restricted to low-dimensional or highly symmetric settings.
In this work we study the Kadison–Schwarz property within the Bloch–Gell–Mann representation of unital maps on $M_{d}$. This parametrization is adapted to the $\mathfrak{su}(d)$ structure and makes it possible to exploit unitary equivalence and canonical forms of the Bloch matrix. While the representation exists in any dimension, obtaining explicit analytic KS criteria is strongly dimension dependent.
We focus on $M_{3}$ and derive an explicit sufficient condition for the KS property for unital positive maps with diagonal Bloch matrix. The proof is based on an explicit expansion of the KS expression in the Gell–Mann basis and on a dominance estimate in which the traceless part is controlled by the spectral spread $\max_{i,j}|\mu_{i}-\mu_{j}|$ of the Bloch parameters. A key structural mechanism is that for diagonal Bloch matrices all contributions involving the antisymmetric structure constants cancel, so that the relevant estimates are governed by the symmetric tensor $d_{ijk}$. We also discuss an illustrative two-parameter diagonal family, which makes the resulting KS region in parameter space transparent and highlights the difference between positivity, the KS property, and complete positivity.
The paper is organized as follows. In Sec. II we recall the definition of Kadison–Schwarz maps and review their basic properties. In Sec. III we introduce the Bloch–Gell–Mann representation of unital maps on $M_{d}$. Our main results, specialized to the case of $M_{3}$, are presented in Sec. IV, followed by an illustrative example. Technical derivations are collected in the Appendix.
@@ II Kadison–Schwarz maps
which is required to hold for all $X\in M_{d}$. The inequality (1) was originally introduced by Kadison in the context of operator algebras [9]. We emphasize that the KS property does not, in general, imply unitality, although the two notions are often considered together.
@@
see, e.g., Refs. [5, 12, 13]. Despite this simple ordering, the structure of KS maps is considerably less understood than that of completely positive maps.
@@
Lemma 1 (Kadison-type contraction).
@@
This estimate follows from Kadison’s inequality
$\Phi(A)^{2}\leq\Phi(A^{2})\qquad(A=A^{\dagger}),$ (4)
which holds for unital positive maps [9, 15, 20]. Indeed, positivity implies that $\Phi$ maps Hermitian elements to Hermitian elements, hence $\Phi(A)=\Phi(A)^{\dagger}$. Therefore $\|\Phi(A)\|^{2}=\|\Phi(A)^{2}\|$, and Kadison’s inequality gives
$\|\Phi(A)\|^{2}=\|\Phi(A)^{2}\|\leq\|\Phi(A^{2})\|.$
Since $A^{2}\geq 0$ and $A^{2}\leq\|A^{2}\|\mathbb{I}$, positivity and unitality imply $0\leq\Phi(A^{2})\leq\|A^{2}\|\Phi(\mathbb{I})=\|A^{2}\|\mathbb{I}$, hence $\|\Phi(A^{2})\|\leq\|A^{2}\|=\|A\|^{2}$.
Lemma 1 implies that unital positive maps are contractions on the self-adjoint part of $M_{d}$ with respect to the operator norm. Importantly, this property does not rely on complete positivity and plays a key role in the analysis of KS maps [5].
It is worth noting that for unital positive maps the Kadison–Schwarz inequality is automatically satisfied when restricted to normal operators, whereas possible violations of the KS condition necessarily involve non-normal elements [5, 13].
@@
Hence
$T(X^{\dagger}X)-T(X^{\dagger})T(X)=|2\rangle\langle 2|-|1\rangle\langle 1|.$
The latter operator has eigenvalues +1+1 and 1-1, and therefore is not positive. In particular, the presence of a negative eigenvalue shows that
$T(X^{\dagger}X)\ngeq T(X^{\dagger})T(X),$
so the Kadison–Schwarz inequality is violated. This example is well known and discussed in detail in the literature, see, e.g., Refs. [5, 13].
@@ III Bloch–Gell–Mann representation of unital maps on $M_{d}$
The Bloch representation of linear maps does not depend on the choice of an orthonormal basis of $\mathfrak{su}(d)$: different choices (e.g., Weyl or generalized Pauli bases) are related by orthogonal transformations in Bloch space and yield equivalent descriptions of unital maps [10, 1]. In particular, the Kadison–Schwarz property is invariant under such basis changes. We use the Gell–Mann basis because its generators are Hermitian and its symmetric and antisymmetric structure constants allow a transparent treatment of products appearing in the KS inequality [7].
@@ III.1 Bloch representation on $M_{d}$
where $w_{0}\in\mathbb{R}$ and $w_{k}\in\mathbb{C}$. For Hermitian operators $X$, all coefficients $w_{k}$ are real [10, 1].
@@
where $f_{ijk}$ and $d_{ijk}$ are the completely antisymmetric and symmetric structure constants of $\mathfrak{su}(d)$, respectively [7, 18].
@@ III.2 Unital maps in the Bloch representation
where $w=(w_{1},\dots,w_{d^{2}-1})^{T}$ and $T$ is a real $(d^{2}-1)\times(d^{2}-1)$ matrix, referred to as the Bloch matrix of $\Phi$ [10, 17]. Unitality of $\Phi$ implies that the identity component $w_{0}$ is preserved, while positivity imposes nontrivial constraints on $T$.
@@
where $O_{1},O_{2}\in SO(d^{2}-1)$ are orthogonal matrices induced by the adjoint action of $SU(d)$ on $\mathfrak{su}(d)$ [1].
@@ III.3 Canonical form
where $R\in O(d^{2}-1)$ is orthogonal and $S$ is symmetric and positive semidefinite. If the orthogonal factor $R$ is induced by unitary pre- and post-processing, i.e., by conjugations $X\mapsto UXU^{\dagger}$ acting on the traceless sector, then the Kadison–Schwarz property is preserved under the corresponding transformation $T\mapsto O_{1}TO_{2}$. For $d\geq 3$, however, not every orthogonal transformation in $SO(d^{2}-1)$ is realized by the adjoint action of $SU(d)$. In what follows we therefore focus on the analytically tractable class of unital maps that are unitarily equivalent (within this physical equivalence) to maps with symmetric, and in particular diagonal, Bloch matrices.
Accordingly, we focus on unital maps whose Bloch matrices can be chosen symmetric within their unitary equivalence class. In particular, if such a symmetric Bloch matrix is diagonalizable, then the map $\Phi$ is unitarily equivalent to a map with diagonal Bloch matrix
@@ IV Main result: Kadison–Schwarz condition on $M_{3}$
Let $\Phi:M_{3}\to M_{3}$ be a unital positive linear map with diagonal Bloch matrix
@@
in the Gell–Mann representation. Then, by Lemma 1, one has
$\max_{1\leq k\leq 8}|\mu_{k}|\leq 1,$ (13)
since $\lambda_{k}=\lambda_{k}^{\dagger}$ and $\Phi(\lambda_{k})=\mu_{k}\lambda_{k}$, hence
$|\mu_{k}|\,\|\lambda_{k}\|=\|\Phi(\lambda_{k})\|\leq\|\lambda_{k}\|$
and therefore $|\mu_{k}|\leq 1$ since $\|\lambda_{k}\|>0$. Define
$c_{3}(\mu):=\inf_{X\neq 0}\frac{\alpha(X)}{\|X^{\dagger}X\|}\ \geq 0,$ (14)
where $\alpha(X)$ is the scalar coefficient in the expansion (16). Note that $\|X^{\dagger}X\|>0$ for $X\neq 0$, so $c_{3}(\mu)$ is well-defined. If
$\max_{i,j}|\mu_{i}-\mu_{j}|\;\leq\;\frac{c_{3}(\mu)}{C_{3}},$ (15)
where $C_{3}$ is the structural constant from Appendix A, then $\Phi$ satisfies the Kadison–Schwarz inequality.
@@
The constant $C_{3}$ controls the traceless contribution in the expansion (16) via the norm estimate derived in Appendix A. The quantity $c_{3}(\mu)$ is defined in (14) and quantifies a uniform lower bound for the scalar term $\alpha(X)$ relative to $\|X^{\dagger}X\|$. Condition (15) is sufficient, but in general not necessary, for the Kadison–Schwarz property within the class of unital positive maps. Note that if $c_{3}(\mu)=0$, then (15) forces $\max_{i,j}|\mu_{i}-\mu_{j}|=0$, so the sufficient criterion becomes nontrivial only when $c_{3}(\mu)>0$. Although $c_{3}(\mu)$ is defined variationally, in concrete subclasses of diagonal maps it can often be bounded from below explicitly; we leave optimization of such bounds for future work.
We stress that the theorem provides a purely analytic sufficient criterion that does not rely on complete positivity or optimization-based arguments, and therefore isolates a genuinely structural mechanism behind the Kadison–Schwarz property on $M_{3}$.
@@ IV.1 Reduction of the Kadison–Schwarz inequality
Lemma 2 (Cancellation of the antisymmetric part for diagonal $T$).
Let $\Phi$ have diagonal Bloch matrix $T=\operatorname{diag}(\mu_{1},\dots,\mu_{8})$ with real parameters $\mu_{k}$. Then in the expansion of the Kadison–Schwarz expression $\Phi(X^{\dagger}X)-\Phi(X^{\dagger})\Phi(X)$ all contributions proportional to the antisymmetric structure constants $f_{ijk}$ cancel.
Proof.
In the Bloch–Gell–Mann product expansion, the $f_{ijk}$-terms appear with coefficients antisymmetric under the exchange $i\leftrightarrow j$, whereas the quadratic combinations $w_{i}\overline{w_{j}}$ arising from $X^{\dagger}X$ are paired with their conjugates $w_{j}\overline{w_{i}}$. This pairing is automatic because the KS expression is Hermitian and the Bloch expansion collects each coefficient together with its complex conjugate counterpart. Since $T$ is diagonal with real $\mu_{k}$, the resulting coefficients are symmetric in $(i,j)$, and hence the total antisymmetric contribution vanishes after summation. ∎
By Lemma 2, for diagonal $T$ the KS expression is governed entirely by the symmetric tensor $d_{ijk}$. (For $d=2$ one has $d_{ijk}\equiv 0$, so this mechanism is absent in the qubit case.)
@@ IV.2 Proof of Theorem 1
In view of the reduction described above, the Kadison–Schwarz condition reduces to the positivity of the operator (16). We estimate separately its scalar and traceless parts.
Scalar contribution. Using the product relations for the Gell–Mann generators one verifies that the coefficient $\alpha(X)$ is a sum of terms of the form
$(1-\mu_{k}^{2})\,|w_{k}|^{2}$
(together with nonnegative contributions proportional to $|w_{0}|^{2}$). Indeed, the identity component in (7) is proportional to $\delta_{ij}$, hence contributes only terms with $i=j$. Hence, under the contraction assumption (13), one has $\alpha(X)\geq 0$ for all $X$.
Moreover, define
$c_{3}(\mu):=\inf_{X\neq 0}\frac{\alpha(X)}{\|X^{\dagger}X\|}\geq 0,$
as in (14). By definition, this implies the uniform bound
$\alpha(X)\ \geq\ c_{3}(\mu)\,\|X^{\dagger}X\|\qquad\text{for all }X\in M_{3}.$ (17)
Traceless contribution. The remaining traceless part can be estimated using operator norm bounds for expressions involving the symmetric structure constants $d_{ijk}$. More precisely, one obtains
@@
where $C_{3}$ depends only on the algebraic structure of $\mathfrak{su}(3)$, as shown in Appendix A.
Dominance argument. If
$\max_{i,j}|\mu_{i}-\mu_{j}|\;\leq\;\frac{c_{3}(\mu)}{C_{3}},$
then combining (17) and (18) yields
$\biggl\|\sum_{k=1}^{8}\beta_{k}(X)\,\lambda_{k}\biggr\|\;\leq\;c_{3}(\mu)\,\|X^{\dagger}X\|\;\leq\;\alpha(X).$
Therefore the traceless part is dominated by the positive scalar part, and the operator (16) is positive. Hence,
@@
for all $X\in M_{3}$, which proves the theorem. A detailed derivation of the bound (18) and an explicit expression for the constant $C_{3}$ are provided in the Appendix. From a structural perspective, the Kadison–Schwarz condition for diagonal Bloch maps on $M_{3}$ can be interpreted as a stability property under controlled spectral anisotropy.
V Illustrative diagonal family
To illustrate the content of Theorem 1, consider a two-parameter family of diagonal Bloch matrices
$T(t,s)=\operatorname{diag}(t,t,t,t,t,t,t,s),$ (19)
with real parameters $t,s\in[-1,1]$. For general $(t,s)$ the associated unital map need not be positive. We do not characterize the positivity region of $T(t,s)$ here and restrict the discussion to parameters for which the map is positive. Such maps act isotropically on a seven-dimensional subspace of $\mathfrak{su}(3)$ while allowing an independent deformation along a distinguished direction.
The contraction condition (13) is satisfied whenever $|t|\leq 1$ and $|s|\leq 1$. Moreover,
$\max_{i,j}|\mu_{i}-\mu_{j}|=|t-s|.$ (20)
Hence Theorem 1 guarantees the Kadison–Schwarz property whenever
$|t-s|\leq\frac{c_{3}(t,s)}{C_{3}},$ (21)
where $C_{3}$ is the structural constant defined in Appendix A and
[…]
This example shows explicitly that the sufficient condition defines a nontrivial region in parameter space, controlled only by the relative spectral spread of the Bloch parameters rather than by their individual values. Geometrically, the admissible region corresponds to a neighborhood of the isotropic line $t=s$ whose width is determined by the ratio $c_{3}(t,s)/C_{3}$, reflecting the balance between the positive scalar part and the traceless fluctuations governed by the $\mathfrak{su}(3)$ structure constants.
Positivity and complete positivity for the isotropic subfamily. For the isotropic choice $t=s$, the map acts as a scalar on the traceless subspace and admits the depolarizing form
$\Phi_{t}(X)=tX+(1-t)\frac{\operatorname{Tr}(X)}{3}\,\mathbb{I}.$ (22)
Applying $\Phi_{t}$ to a rank-one projection $P_{\psi}=|\psi\rangle\langle\psi|$ yields
$\Phi_{t}(P_{\psi})=tP_{\psi}+\frac{1-t}{3}\mathbb{I},$
whose eigenvalues are $\lambda_{\parallel}=(2t+1)/3$ and $\lambda_{\perp}=(1-t)/3$ (with multiplicity $2$). Hence the map is positive if and only if
$\frac{1}{2}\leq t\leq 1,$ (23)
see [21, 14]. The corresponding Choi matrix $J(\Phi_{t})=(\Phi_{t}\otimes\operatorname{id})(|\Phi\rangle\langle\Phi|)$, with $|\Phi\rangle=\frac{1}{\sqrt{3}}\sum_{i=1}^{3}|ii\rangle$, takes the form
$J(\Phi_{t})=t\,|\Phi\rangle\langle\Phi|+\frac{1-t}{9}\,\mathbb{I}\otimes\mathbb{I}.$ (24)
Its spectrum consists of the eigenvalue
$\lambda_{1}=\frac{8t+1}{9}$
and the eigenvalue
$\lambda_{2}=\frac{1-t}{9}$
with multiplicity $8$. Therefore $\Phi_{t}$ is completely positive if and only if
$\frac{1}{8}\leq t\leq 1,$ (25)
see [21, 14, 8].